OpenStudy (megannicole51):

Use Trig Substitution to solve the integral of ((dt)/(t^2)(square root of 1+t^2))

OpenStudy (anonymous):

no more partial fractions?

OpenStudy (megannicole51):

nope! im good with partial fractions! i can do those in my sleep:)

OpenStudy (megannicole51):

my professor taught Trig Substitution with just letters....no numbers in the examples or anything so I am so confused its crazy

OpenStudy (anonymous):

okay... just a sec.

OpenStudy (megannicole51):

thank you:)

OpenStudy (anonymous):

lol partial fractions for days

OpenStudy (anonymous):

@pgpilot326 can probably do this better, but if i remember correctly when you see \[\sqrt{1+x^2}\] you use \(x=\tan(\theta)\)

OpenStudy (anonymous):

because \[1+\tan^2(\theta)=\sec^2(\theta)\]

OpenStudy (megannicole51):

yup ive got that far! lol

OpenStudy (anonymous):

so now if x = tan theta, what's dx?

OpenStudy (megannicole51):

sec^2(theta)

OpenStudy (megannicole51):

theres a d(theta) after that sorry

OpenStudy (anonymous):

do some trig simplification there will be a ton of cancellation

OpenStudy (anonymous):

\[\frac{\sec^2(\theta)}{\tan^2(\theta)sec(\theta)}\] is the first step, but tons goes

OpenStudy (anonymous):

so then this becomes \[\int\limits\frac{ dt }{t^{2}\sqrt{1+t^{2} }}= \int\limits\frac{ \sec ^{2} \theta d \theta }{\tan ^{2} \theta \sec \theta }\]

OpenStudy (megannicole51):

okay :) just pluggin in easy enough

OpenStudy (anonymous):

hint: you end up with something you know the anti derivative of instantly (i think)

OpenStudy (anonymous):

yep... just like @satellite73 said, there's cancellation. don't forget to draw the triangle used for the trig sub as you may need other things from it after integrating

OpenStudy (megannicole51):

yup ive got the triangle! :)

OpenStudy (megannicole51):

so am i able to cancel out the secants with one left in the numerator? or can i not do that?

OpenStudy (anonymous):

yep

OpenStudy (megannicole51):

so am i just left with -csc(theta)+c

OpenStudy (anonymous):

yes, but now back to x

OpenStudy (megannicole51):

im not sure how to go back

OpenStudy (anonymous):

use the triangle

OpenStudy (megannicole51):

i know how to write out the triangle but im not real sure what to do with it:(

OpenStudy (megannicole51):

sorry:(

OpenStudy (anonymous):

draw your triangle and i'll respond and show you

OpenStudy (anonymous):

i bet it will be clear once you get it

OpenStudy (anonymous):

csc = h/o

OpenStudy (megannicole51):